OpenStudy (anonymous):

Mary takes 3 hours to mow the yard. It takes Sally 5 hours. How long would it take with them mowing it together? Give the exact decimal answer. A. 3.4 hours B. 2.725 hours C. 1.875 hours D. 2.225 hours

OpenStudy (anonymous):

In one hour they both can do 1/x 1/3 + 1/5 = 1/x 8/15 = 1/x x = 1.875 hours Hope this helps!